# references.bib

### v2

@@ IV-C Zero-shot Relative Depth Estimation
[…]
@@
AbsRel↓ SqRel↓ RMSE↓ RMSElog↓ $\delta 1$ AbsRel↓ SqRel↓ RMSE↓ RMSE log↓ $\delta 1$
Depth Anything [yang2024depth] 0.041 0.25 3.986 0.053 0.996 0.100 1.330 7.858 0.126 0.869
Depth Anything v2 [yang2024depth2] 0.040 0.251 3.797 0.053 0.994 0.099 1.301 7.815 0.125 0.872
SSI [ranftl2020towards] 0.038 0.174 3.420 0.047 0.999 0.098 1.304 7.756 0.125 0.874
TSSI [ranftl2020towards] 0.037 0.169 3.399 0.047 1.000 0.100 1.331 7.850 0.126 0.870
Ours: WSSI-labeled 0.037 0.174 3.418 0.046 1.000 0.098 1.294 7.725 0.124 0.873
Ours: WSSI 0.034 0.158 3.316 0.044 1.000 0.097 1.286 7.687 0.123 0.875
@@
TABLE VIII: Fine-tuned to segmentation on Kvasir dataset.
Methods Dice mIoU Recall Precision F2
@@
Ours 0.909 0.857 0.912 0.939 0.909
TABLE IX: Branch-level endoscope localization accuracy by fine-tuning to airway lumen detection.
@@ IV-D Fine-tuned Metric Depth Estimation
Estimating depth with an absolute scale is critical for downstream tasks such as localization and mapping. However, only a limited amount of existing work in medical imaging focuses on predicting metric depth from monocular endoscopic images, likely due to the relatively small size of each individual dataset. Thus, we aim to provide a powerful foundation model as a promising initialization for fine-tuning MDE using a limited amount of data. Specifically, we initialize the encoder weight with our pretrained student model $f$ and randomly initialize the decoder. The network is fine-tuned on a single dataset following the training strategy by [bhat2023zoedepth].
@@ IV-E Ablation Studies
Robust Training Loss. We validate the effectiveness of our proposed robust training loss by analyzing the zero-shot RDE performance following the first two stages of training. Since the second stage employs a scale-and-shift invariant loss, we evaluate RDE by aligning the per-frame scale and shift with ground truth. We compare the performance of four different training losses: (1) SSI, the standard scale-and-shift invariant loss initially proposed by MiDaS [ranftl2020towards]; (2) TSSI, a loss proposed by [ranftl2020towards], where outliers are identified and removed by trimming the largest loss residuals per instance; (3) Ours: WSSI-labeled, which incorporates the teacher model to assess ground truth confidence, enhancing learning bias for labeled data; and (4) Ours: WSSI, which is our complete weighted scale and shift invariant loss that extends confidence analysis to both labeled and unlabeled data. We evaluate these losses on the SERV-CT and Hamlyn datasets. As shown in Table VII, our WSSI consistently outperforms other methods across all metrics. Interestingly, we observe that trimming loss outliers, while intended to detect label noise, does not perform as expected on medical images. This supports our hypothesis that the trimming process may also discard hard-to-learn regions of each instance.
Additionally, we examine the training dynamics of the different losses, as shown in Fig. 6. Both WSSI and TSSI contribute to training stability, while the standard SSI loss leads to training collapse due to label noise interference. This results in premature underfitting, with the best-performing model on the validation set producing depth maps with blurred edges, as shown in Fig. 5. The performance degradation is particularly evident on the SERV-CT dataset, where accurate dense depth labels highlight the deficiencies of the vanilla SSI loss compared to the Hamlyn dataset.
Comparison with Depth Anything. Since EndoOmni initializes encoder weights from the general foundation model Depth Anything, we compare our zero-shot RDE performance to demonstrate the improvements on endoscopy data. We also include a comparison with the latest Depth Anything v2 for reference. As shown in Table VII, our EndoOmni outperforms all baselines across the evaluated metrics, particularly on the SERV-CT dataset, which offers more accurate dense depth labels, highlighting our model’s superior performance.
@@ IV-F Fine-tuned to Polyp Segmentation
Although originally trained for the SDE task, EndoOmni is expected to transfer effectively to other downstream tasks in endoscopy. To validate this capability, we applied EndoOmni to the task of polyp segmentation. We followed the data split of the established benchmark [fan2020pranet] and conducted experiments on the Kvasir-SEG dataset [jha2020kvasir]. For this task, we modified EndoOmni by removing the final ReLU activation and adding batch normalization and dropout layers, without introducing any task-specific modules for segmentation. The results, shown in Table VIII, demonstrate that EndoOmni achieves superior performance compared to existing methods in medical imaging segmentation, highlighting its strong semantic representation capabilities. Additionally, compared with initialization with DINOv2 [oquab2024dinov2], we observed a significant improvement in Dice score from 0.883 to 0.909, underscoring EndoOmni’s potential as a versatile foundation model for multi-task applications in endoscopy.
@@ IV-G Fine-tuned to Bronchoscopy Localization
To further evaluate EndoOmni’s transferability to other endoscopic tasks, we fine-tuned it to airway lumen detection. This problem, initially introduced by [sganga2019autonomous] and later used for branch-level endoscope localization in the airway [tian2024bronchotrack]. We hypothesize that the geometric structure semantics learned during SDE training enable EndoOmni to transfer learning to anatomical understanding effectively. We initialized the pretrained EndoOmni encoder and employed a 3-layer MLP to detect airway branches in each frame. The final endoscope localization accuracy was evaluated using the same test set from [tian2024bronchotrack]. As shown in Table IX, the fine-tuned EndoOmni outperformed existing specialized methods for branch-level endoscope localization. Additionally, we tested initializing the encoder with the powerful DINOv2 [oquab2024dinov2] pretrained weights, and EndoOmni still outperformed with a comfortable margin. This further indicates that EndoOmni could serve as a robust foundation for other downstream tasks in endoscopy.
V Conclusion

### v3

@@ IV-C Zero-shot Relative Depth Estimation
Compared with existing research on SDE in medical imaging, one important advantage of leveraging a foundation model trained on abundant mixed data is accurate depth estimation on unseen datasets during training. Therefore, we validate zero-shot RDE performance on two endoscopic datasets: Hamlyn and SERV-CT [edwards2020serv]. We follow existing works on MDE in medical imaging, fine-tuning the model on the SCARED dataset [allan2021stereo] at the third stage of training of our foundation model and use the same split scheme as existing works [recasens2021, cui2024surgical, cui2024endodac]. As shown in Table II and Table III, our model surpasses existing methods by a large margin. Notably, the best performing existing state-of-the-art (SOTA) method on the Hamlyn and SERV-CT Dataset, EndoDAC [cui2024endodac], focuses on efficiently adapting Depth Anything-B [yang2024depth] to endoscopic scenes through parameter-efficient fine-tuning. Compared with EndoDAC, our EndoOmni-B improves on all evaluated metrics by a large margin. These significant improvements highlight the advantages of robust learning with large, diverse data, which supports training larger models and learning more generalizable features. Our model thus gains a specialized feature understanding of medical data through our curated endoscopy datasets and advanced robust training framework. Quantitative results are shown in Fig. 5. Our EndoOmni provides the highest-quality outputs with well-defined edges.
@@
Abs Rel↓ Sq Rel↓ RMSE↓ RMSE log↓ $\delta$1↑ Abs Rel↓ Sq Rel↓ RMSE↓ RMSE log↓ $\delta$1↑
Depth Anything 0.050 0.424 5.590 0.065 0.987 0.111 2.025 9.843 0.140 0.852
Depth Anything v2 0.049 0.422 5.556 0.066 0.987 0.110 1.977 9.847 0.140 0.852
SSI 0.044 0.332 5.115 0.057 0.993 0.110 1.996 9.761 0.139 0.854
TSSI 0.044 0.320 5.074 0.058 0.995 0.109 1.970 9.705 0.139 0.856
Ours: WSSI-labeled 0.042 0.295 4.853 0.054 0.995 0.108 1.936 9.569 0.137 0.857
Ours: WSSI 0.040 0.270 4.716 0.053 0.997 0.106 1.830 9.362 0.135 0.864
@@
TABLE VIII: Ablation on the size of network backbone.
Backbone SERV-CT Hamlyn
AbsRel↓ SqRel↓ RMSE↓ RMSElog↓ $\delta$1↑ AbsRel↓ SqRel↓ RMSE↓ RMSElog↓ $\delta$1↑
ViT-S 0.048 0.357 5.397 0.061 0.995 0.110 1.827 9.843 0.141 0.855
ViT-B 0.044 0.311 4.982 0.058 0.994 0.104 1.653 9.325 0.134 0.872
ViT-L 0.040 0.270 4.716 0.053 0.997 0.106 1.830 9.362 0.135 0.864
TABLE IX: Fine-tuned to segmentation on Kvasir dataset.
Methods Mean Dice mIoU Recall Precision F2
@@
EndoOmni (Ours) 0.909 0.857 0.912 0.939 0.909
TABLE X: Branch-level endoscope localization accuracy by fine-tuning to airway lumen detection.
@@ IV-D Fine-tuned Metric Depth Estimation
Estimating depth with an absolute scale is critical for downstream tasks such as localization and mapping. However, only a limited amount of existing work in medical imaging focuses on predicting metric depth from monocular endoscopic images, likely due to the relatively small size of each individual dataset. Thus, we aim to provide a powerful foundation model as a promising initialization for fine-tuning MDE using a limited amount of data. Specifically, we initialize the encoder weight with our pretrained student model $f$ and randomly initialize the decoder. The network is fine-tuned on a single dataset following the training strategy by [bhat2023zoedepth]. We used EndoOmni with the ViT-L backbone for the fine-tuning to metric depth experiments, following [yang2024depth].
@@ IV-E Ablation Studies
Robust Training Loss. We validate the effectiveness of our proposed robust training loss by analyzing the zero-shot RDE performance following the first two stages of training. Since the second stage employs a scale-and-shift invariant loss, we evaluate RDE by aligning the per-frame scale and shift with ground truth. We compare the performance of four different training losses: (1) SSI, the standard scale-and-shift invariant loss initially proposed by MiDaS [ranftl2020towards]; (2) TSSI, a loss proposed by [ranftl2020towards], where outliers are identified and removed by trimming the largest loss residuals per instance; (3) Ours: WSSI-labeled, which incorporates the teacher model to assess ground truth confidence, enhancing learning bias for labeled data; and (4) Ours: WSSI, which is our complete weighted scale and shift invariant loss that extends confidence analysis to both labeled and unlabeled data. We evaluate these losses on the SERV-CT and Hamlyn datasets. As shown in Table VII, our WSSI consistently outperforms other methods across all metrics on both datasets, demonstrating the robustness of WSSI in handling diverse data distributions. Interestingly, while TSSI shows minor improvements over SSI, the performance gain is marginal when compared to WSSI. This reinforces the idea that trimming loss outliers is not the most efficient approach for improving performance on medical image datasets.
Additionally, we examine the training dynamics of the different losses, as shown in Fig. 6. Both WSSI and TSSI contribute to training stability, while the standard SSI loss leads to training collapse due to label noise interference. This results in premature underfitting, with the best-performing model on the validation set producing depth maps with blurred edges, as shown in Fig. 5.
Comparison with Depth Anything. Since EndoOmni initializes encoder weights from the general foundation model Depth Anything, we compare our zero-shot RDE performance to demonstrate the improvements in endoscopy data. We also include a comparison with the latest Depth Anything v2 for reference. As shown in Table VII, our EndoOmni outperforms all baselines across the evaluated metrics, particularly on the SERV-CT dataset, which offers more accurate dense depth labels, highlighting our model’s superior performance.
Ablation on Backbone Size. To accommodate varying application requirements, we provide three student model sizes based on the ViT-S (24.8M parameters), ViT-B (97.5M parameters), and ViT-L (335.3M parameters) architectures. Their zero-shot performance results are presented in Table VIII. The results indicate that EndoOmni’s performance shows only a slight decline as the backbone parameter count decreases, making it suitable for resource-constrained environments.
@@ IV-F Fine-tuned to Polyp Segmentation
Although originally trained for the SDE task, EndoOmni is expected to transfer effectively to other downstream tasks in endoscopy. To validate this capability, we applied EndoOmni with the ViT-L backbone to the task of polyp segmentation. We followed the data split of the established benchmark [fan2020pranet] and conducted experiments on the Kvasir-SEG dataset [jha2020kvasir]. For this task, we modified EndoOmni by removing the final ReLU activation and adding batch normalization and dropout layers, without introducing any task-specific modules for segmentation. The results, shown in Table IX, demonstrate that EndoOmni achieves superior performance compared to existing methods in medical imaging segmentation, highlighting its strong semantic representation capabilities. Notably, we observed a significant improvement in Dice score from 0.883 to 0.909. Additionally, our EndoOmni adaptation for the segmentation task achieved compatible performance with Polyp-SAM-L [li2024polyp], a fine-tuned Segment Anything (SAM) [kirillov2023segment] foundation model. This observation underscores EndoOmni’s potential as a versatile foundation model for multi-task applications in endoscopy.
@@ IV-G Fine-tuned to Bronchoscopy Localization
To further evaluate EndoOmni’s transferability to other endoscopic tasks, we fine-tuned it to airway lumen detection. This problem was initially introduced by [sganga2019autonomous] and later used for branch-level endoscope localization in the airway [tian2024bronchotrack]. We hypothesize that the geometric structure semantics learned during SDE training enable EndoOmni to transfer learning to anatomical understanding effectively. We initialized the pretrained EndoOmni encoder with the ViT-L backbone and employed a 3-layer MLP to detect airway branches in each frame. The final endoscope localization accuracy was evaluated using the same test set from [tian2024bronchotrack]. As shown in Table X, the fine-tuned EndoOmni outperformed existing specialized methods for branch-level endoscope localization. Additionally, we tested initializing the encoder with the powerful DINOv2 [oquab2024dinov2] pretrained weights, and EndoOmni still outperformed with a comfortable margin. This further indicates that EndoOmni could serve as a robust foundation for other downstream tasks in endoscopy.
V Discussion
Effective generalization to new scenes is crucial for depth estimation methods in real-world applications, especially in medical imaging, where data collection and model fine-tuning are particularly challenging in clinical settings. However, most existing research in medical depth estimation has focused on specific datasets [recasens2021, shao2022self, ozyoruk2021endoslam, puigvert2023lightdepth, li2024image, wang2024monopcc, tian2024dd, mathew2020augmenting, shen2019context, wei2024absolute].
In contrast, foundation models in general computer vision excel in zero-shot SDE and offer robust generalization across diverse scenes. However, models trained on general scenes suffer considerable performance drops when applied to specialized domains like endoscopic imaging [cui2024endodac]. Consequently, recent efforts have adapted foundation models to medical datasets through parameter-efficient fine-tuning [cui2024surgical, cui2024endodac], with some focusing on in-domain testing [li2024advancing]. These adaptations are typically a necessary compromise due to lacking an off-the-shelf solution for depth estimation in medical contexts. However, this fine-tuning can lead to representational collapse, which may reduce the foundation models’ generalization capabilities across broader domains [aghajanyan2020better].
EndoOmni directly addresses this gap by offering a robust, off-the-shelf solution for endoscopic SDE that performs well even without fine-tuning. Our approach builds on the straightforward yet effective method of creating a comprehensive endoscopic meta-dataset to train our foundation model. By leveraging the largest meta-dataset for medical SDE, sourced from both public and proprietary datasets, and applying the teacher-student paradigm [hinton2015distilling], we maximize the use of unlabeled data. Ablation studies (SSI results) show that this approach yields substantial improvements over general-domain foundation models.
Another key contribution of this work is addressing the training instability observed when using the standard SSI loss [ranftl2020towards] on medical data, resulting in depth predictions often exhibiting blurry edges (see Fig. 5, SSI column). In contrast, general-domain foundation models [yang2024depth] produce sharper edges (see Fig. 5, Depth Anything column). To tackle this, we developed a refined self-learning framework with a robust WSSI loss. Our analysis shows that clean labels converge more reliably, and pseudo-labels become more dependable when the teacher model’s predictions are consistent across augmentations. Building on this, we estimate per-pixel confidence from the teacher model’s behavior, guiding student model training with both labels and confidence. The WSSI loss dynamically adjusts learning weights based on confidence, mitigating noise while preserving challenging regions, ultimately enhancing depth estimation performance.
The TSSI loss proposed by MiDas [ranftl2020towards], by contrast, employs a trimming strategy that excludes high-loss components, which can overlook complex image regions. Our ablation studies indicate that TSSI does not consistently outperform SSI in medical imaging, as evidenced by results in Table VII. WSSI reliably outperforms both SSI and TSSI, achieving statistically significant gains on all metrics with the Hamlyn dataset.
While we curated the largest meta-dataset for endoscopic depth estimation, additional datasets could further improve generalizability. Future work will expand our dataset with more labeled and unlabeled data for training. Although this study emphasizes robustness against label noise, endoscopic data also presents unique challenges from image artifacts like occlusions, bubbles, and specular reflections. We aim to address these issues in future work by developing robust learning methods tailored to the complexities of endoscopic imaging. Additionally, while we provide a fine-tuning approach for metric depth estimation in endoscopic imaging—vital for tasks like localization and navigation [tian2024pans]—achieving zero-shot metric depth estimation across varied camera systems remains an open challenge [bhat2023zoedepth, piccinelli2024unidepth], which we leave for future exploration.
VI Conclusion